# HACTS: a Human-As-Copilot Teleoperation System for Robot Learning

### v1

@@ IV-E Experimental Results on Reinforcement Learning
We developed HACTS-RLPD by enhancing existing reinforcement learning methods with a human-in-the-loop, vision-based system. In this setup, a robot arm autonomously explores its environment but receives human assistance during challenging situations or when guidance is needed. This human input is seamlessly integrated into the learning process, enabling the RL algorithms to refine the control policy progressively by utilizing data from both robot and human-copilot actions.
Answer for Question 4. As shown in Tab. IV, HACTS-RLPD achieves 80% success rates after 10 minutes offline pretraining and 45 minutes online training. Compared with BC that trained only with human demonstrations, the policy trained by RL demonstrate shorter episode length due to the human assisted trail-and-error learning manner and the sparse reward settings.
TABLE IV: Results of HACTS-RLPD from offline BC pretraining followed by online finetuning. We report the success rates and episode lengths over 10 trials.
@@
Figure 6: Average intervention length during HACTS-RLPD training in the CloseBin Task. Each episode terminates either after exceeding 200 steps or when the reward classifier outputs a logit value exceeding 0.85.
In Figure 6, we present the average human intervention length during online training, computed as the total number of intervention steps divided by the times of consecutive interventions. Initially, intervention length increases, reflecting a performance drop in HACTS-RLPD due to Q-network learning and exploration. During this phase, humans must provide long interventions to ensure success. As training progresses, the policy increasingly relies on its Q estimations, reducing the need for human input. Eventually, interventions drop below six steps, enabling fine-grained corrections and preventing overly aggressive closing strategies near the end of each episode.
@@ V Conclusion and Discussion
Through the implementation of low-cost, kinematically equivalent components, we demonstrate that HACTS offers a scalable and accessible solution for teleoperation, making it a promising tool for advanced robot learning applications. Our experiments showed that integrating action-correction data into training significantly enhances the performance of Vision-Language-Action (VLA) models, improving their recovery and data efficiency. Furthermore, the system supports more complex human-in-the-loop reinforcement learning, enabling robots to adapt more effectively in dynamic environments.

### v2

@@ IV-E Experimental Results on Reinforcement Learning
We developed RLPD-HACTS by enhancing existing reinforcement learning methods with a human-in-the-loop, vision-based system. In this setup, a robot arm autonomously explores its environment but receives human assistance during challenging situations or when guidance is needed. This human input is seamlessly integrated into the learning process, enabling the RL algorithms to refine the control policy progressively by utilizing data from both robot and human-copilot actions.
Answer for Question 4. As shown in Tab. IV, RLPD-HACTS achieves 80% success rates after 10 minutes offline pretraining and 45 minutes online training. Compared with BC that trained only with human demonstrations, the policy trained by RL demonstrate shorter episode length due to the human assisted trial-and-error learning manner and the sparse reward settings.
TABLE IV: Results of RLPD-HACTS from offline BC pretraining followed by online finetuning. We report the success rates and episode lengths over 10 trials.
@@
Figure 6: Average intervention length during RLPD-HACTS training in the CloseBin Task. Each episode terminates either after exceeding 200 steps or when the reward classifier outputs a logit value exceeding 0.85.
In Fig. 6, we present the average human intervention length during online training, computed as the total number of intervention steps divided by the times of consecutive interventions. Initially, intervention length increases, reflecting a performance drop in RLPD-HACTS due to Q-network learning and exploration. During this phase, humans must provide long interventions to ensure success. As training progresses, the policy increasingly relies on its Q estimations, reducing the need for human input. Eventually, interventions drop below six steps, enabling fine-grained corrections and preventing overly aggressive closing strategies near the end of each episode.
@@ V Conclusion and Discussion
Through the implementation of low-cost, kinematically equivalent components, we demonstrate that HACTS offers a scalable and accessible solution for teleoperation, making it a promising tool for advanced robot learning applications. Our experiments showed that integrating action-correction data into training significantly enhances the performance of visuomotor policies, improving their recovery and data efficiency. Furthermore, the system supports more complex human-in-the-loop reinforcement learning, enabling robots to adapt more effectively in dynamic environments.